# Multi-person eye tracking for real-world scene perception in social settings

### v1

@@ Abstract.
Eye movements provide a window into human behaviour, attention, and interaction dynamics. Previous research suggests that eye movements are highly influenced by task, setting, and social others; however, most eye tracking research is conducted in single-person, in-lab settings and is yet to be validated in multi-person, naturalistic contexts. One such prevalent real-world context is the collective viewing of a shared scene in social settings, for example, viewing a concert, film, lecture, sporting event, etc. Here, we apply mobile eye tracking in a real-world multi-person setup and develop a system to stream, record, and analyse synchronised data. We tested our proposed, open-source system while participants (N=60) watched a live concert and a documentary film screening during a public event. We tackled challenges related to networking bandwidth requirements, real-time monitoring, and gaze projection from individual egocentric perspectives to a common coordinate space for shared gaze analysis. Our system achieves precise time synchronisation and accurate gaze projection in challenging dynamic scenes. Further, to illustrate the potential of collective eye-tracking data, we introduce and evaluate novel analysis metrics and visualisations. Overall, our approach contributes to the development and application of versatile multi-person eye tracking systems in real-world social settings. This advancement enables insight into collaborative behaviour, group dynamics, and social interaction, with high ecological validity. Moreover, it paves the path for innovative, interactive tools that promote collaboration and coordination in social contexts.
multi-person, eye tracking, social settings, naturalistic, unrestricted, analysis, visualisation, gaze, collective viewing, music, concert, film, homography, modular, system design
$c$s: Human-centered computing Empirical studies in collaborative and social computingccs: Human-centered computing Computer supported cooperative workccs: Human-centered computing Collaborative and social computing design and evaluation methodsccs: Human-centered computing Visualization toolkits
@@ 1. Introduction
The eyes can reveal where, when, and to what we are paying attention (Klein and Ettinger, 2019). Eye contact and shared gaze are critical in social interactions and in guiding collective/joint attention (Valtakari et al., 2021; Clark, 1985). Though ocular activity has always been considered relevant in visual situations, the eyes are equally relevant in indexing auditory attention (Reisberg et al., 1981; Van der Stigchel and Hollingworth, 2018; Pomper and Chait, 2017; Fink et al., 2018, 2019). Interactions between auditory and visual stimuli can capture attention (Van der Burg et al., 2008), alter (Sekuler et al., 1997; Dunifon et al., 2016) or enhance (Bolger et al., 2014; Miller et al., 2013) perceptual sensitivity, and affect social judgments and feelings (Gerdes et al., 2021; Stupacher et al., 2020). Eye tracking can, therefore, provide a multimodal index to perception, attention, and subjective states.
@@
Applications of eye tracking have been mostly limited to single-person settings in controlled environments due to the immobility of traditional eye tracking hardware. With the recent development of affordable, mobile eye tracking technology, it is increasingly possible to conduct eye tracking studies “in the wild” (see e.g. (Fasold et al., 2021; Kulke and Hinrichs, 2021; Reitstätter et al., 2020; Saxena et al., 2023; Vandemoortele et al., 2018; Fridman et al., 2018)). The ability to study unrestricted eye movement behaviour in naturalistic contexts is critical for understanding how eye movements are integrated with head and body movements (Land, 2009) and how previous laboratory results may differ in natural environments (Kulke and Hinrichs, 2021; Laidlaw et al., 2016; Foulsham and Kingstone, 2017). In single-person use cases, mobile eye tracking glasses are commonly applied; however, since these glasses record individual gaze from the egocentric perspective of the wearer, they introduce a new challenge of aligning data from differing perspectives when conducting multi-person eye tracking studies. To effectively study collective gaze dynamics in social settings, it is crucial to develop automated ways to synchronise and analyse the eye tracking data of multiple individuals in unrestricted real-world contexts.
@@
One particularly interesting and ubiquitous context is that of collective viewing of a shared scene in social settings where multiple people share a common gaze goal, e.g., sporting events (Kredel et al., 2017), surgery operations (Tolvanen et al., 2022), school classrooms (Jarodzka et al., 2021), etc. Eye tracking in such naturalistic contexts is crucial for studying social experiences that unfold over longer timescales. For instance, in the case of music, while we have learned much via laboratory studies with single participants and short sound clips presented in the lab, there is a push in the field for studying musical engagement in naturalistic social contexts (Tervaniemi, 2023), as the types of experiences people report as deeply meaningful often involve social settings with music that evolves over long periods of time. However, to date, very few solutions exist for multi-person eye tracking in such dynamic real-world settings.
A schematic of our general approach to solving the problem of multi-person eye-tracking in social settings, and automating data analysis from multiple participant perspectives, is detailed in Figure 1. We imagine a scenario where multiple people observe a shared scene from different viewing angles. The eye movements of each individual viewer can be recorded independently using mobile eye tracking glasses. However, to study collective gaze patterns, the recorded wearer-specific data needs to be synchronised and mapped onto a shared coordinate space (semantic gaze mapping). We record a central view of the shared gaze area from a camera placed behind all observers. We employ Network Time Protocol (NTP) to temporally synchronise data from all devices. The synchronised eye-tracking data is then mapped to the unified central coordinate space via feature mapping between the recorded images captured from each wearer’s head-mounted camera and the images from the central camera. Applying a pin-hole camera model, the scenes captured from each wearer’s perspective can be assumed to lie on a planar surface at an infinite distance. The projective transformation between any two views of the same scene is then mathematically related by a planar homography. Because dynamic real-world social settings likely involve challenging artefacts caused by illumination changes, motion distortion, and low-textured regions, and occlusions, we apply a deep learning based method to robustly detect and match feature points. Paired feature points are used to compute a homography matrix that relates each wearer’s view to a central view of the shared scene. Every individuals’ gaze can then be projected onto the shared centralview for subsequent analyses.
To realise this approach, we first design a scalable system architecture for multi-person mobile eye tracking in social settings, agnostic to hardware requirements. We then test our proposed system during two public events in which we record eye data from 30 participants simultaneously (N=60). In validating our approach, we provide visualisation and analysis metrics for multi-person eye tracking data. Overall, we make the following contributions in this paper:
@@
    Develop and validate a scalable multi-person eye tracking system, with flexible modes of operation. Our system provides users the means to monitor incoming data in real-time and to remotely control all associated mobile eye tracking devices in parallel. We tested our system with 30 mobile eye tracking devices in parallel, limited only by the number of devices we own. Our system should easily handle more devices when operating in recording mode. This tool opens unprecedented avenues for batch data collection and real-time interactivity in social settings.
@@
    Solve the problem of multi-perspective gaze data, in the context of shared social scenes. Using our proposed homography approach, it is possible to project the gaze of all eye tracking glasses wearers from egocentric coordinates to shared central coordinates. This technique revolutionises possibilities for automating analyses of social gaze dynamics in real-world settings.
@@
    Introduce and implement standard and new visualisation and analysis techniques for multi-person eye tracking data. We provide solutions for visualising both egocentric and gaze-transformed data and heatmaps, with methods to visualise all individually or in parallel. The gaze-transformed heatmaps offer particular promise in analysing group attention and perception. Similarly, analysis metrics like convex hull area may offer novel insights into collective gaze behaviour.
Figure 1. LEFT: Schematic of a shared scene depicting multiple people wearing mobile eye-tracking glasses gazing at a stage. A central camera at the back of the audience records the shared scene (centralview). MIDDLE: Egocentric worldviews of each glasses wearer need to be projected onto the centralview. Point correspondences between the egocentric and centralviews are used to calculate homography matrices that relate the two views. RIGHT: Using the computed homography matrices to remap the gaze coordinates, all individuals’ gazes can be projected onto the shared scene (centralview).
@@ 2. Related Work
Previous work on multi-person eye tracking has relied on smaller groups (predominantly dyads) and elaborate experimental setups that hinder natural behaviour (Bise et al., 2024; Fasold et al., 2021; Macdonald and Tatler, 2018; Kera et al., 2016; Jermann et al., 2012; Kawase, 2009; Yamada et al., 2014). Most studies present static scenes on a screen as a proxy to predict real-world eye tracking behaviour which was identified as an inefficient approach by Foulsham and Kingstone (2017). Some studies have applied head pose as a proxy for gaze in multi-person setups (Stiefelhagen, 2002; Beyan et al., 2018); however, existing research also shows that head orientation is not always correlated with gaze in social-group interactions (Vrzakova et al., 2016).
@@
In addition to the experimental setup, analysis of multi-person eye tracking data remains a significant hurdle in conducting large scale studies. Previous studies have heavily relied on manual techniques like incorporating gaze reports in surveys (Pennill and Timmers, 2017), mapping gaze to visual stimuli using proprietary closed-source manufacturer software, or hand-coding annotation of region-of-interest (Benjamins et al., 2018; Geeves et al., 2014). These techniques are incredibly time-consuming, with poor scalability to large datasets. Further, they are not entirely reproducible, and are prone to human errors. A common strategy applied to simplify the problem is to calculate coarse gaze measures, such as discrete binary classification of gaze on regions of interest (Müller et al., 2018; Fasold et al., 2021). While more scalable, such methods provide limited-resolution data, thereby reducing the cost-effectiveness of experiments and limiting the insights that can be gained.
@@
More scalable and automated attempts at unrestricted eye tracking are actively being made under the umbrella field of ‘pervasive eye tracking’ (Kasneci, 2017) and motivated by advances in portable head-mounted eye trackers that have recently become more affordable and accurate. Further, the increased availability of open-sourced and automated analysis tools (Panetta et al., 2019; Li et al., 2006) has aided the adoption and potential applications of these mobile eye trackers, which are increasingly being employed in unrestricted natural environments with static scenes, such as art galleries (Reitstätter et al., 2020). However, to date, a synchronised multi-person eye tracking study in dynamic real world social settings has not been done and consequently the analysis methods to evaluate such large-scale data are not readily available. To solve these challenges we conceptualise and implement a system framework for multi-person eye tracking, described in the next section.
3. System Design
At the highest abstraction level, reliable multi-person eye-tracking relies on four critical components: a) the hardware required to collect eye tracking measures and transmit them to a central server over the network with low latency b) the software that efficiently manages multiple devices and synchronises received data from the devices in a fault-tolerant manner, c) storage to efficiently store raw and/or processed data, and d) analysis and visualisation methods. Proper choice of hardware and storage solutions is largely dependent on case-specific infrastructure limitations but it is critical to avoid performance bottlenecks. Here, we conceptualise the software framework, agnostic to specific hardware and storage solutions. We then adapt the proposed framework in a real-world utility study, discuss our hardware and storage choices, and demonstrate analysis and visualisations for multi-person eye-tracking data in collective, social settings.
@@ 3. System Design
In our proposed software framework, presented in Figure 2, incoming data consists of three parallel streams: i) gaze coordinates from each eye tracking device ii) worldview video recorded from a head-mounted camera on each device, and iii) a centralview video of the shared scene, captured and streamed by a stationary camera mounted to the centre of the scene. Incoming data streams are handled by separate modules that engage based on the mode of operation. Modules in our framework abstract information and allow isolated development and updates of separate components.
@@ 3.1.1. Recording
The recording mode is meant for offline operation where data are collected and stored for post-hoc analysis in a separate stage. Such operation is desirable when real-time processing of incoming data is not required, e.g., in research and user studies. Splitting the two stages reduces network bandwidth and computation requirements, thereby allowing complex analysis and long-duration recordings with limited resources.
@@ 3.1.2. Streaming
The streaming operation is required for real-time or online applications where data are collected and processed concurrently. This mode of operation demands sophisticated hardware and software optimisations to achieve good performance. Processing data in real-time is desirable for innovative Human-Computer Interaction applications such as gaze-contingent interfaces.
Apache Kafka serves as the backbone of the online mode of operation. It is a high-throughput, low-latency event streaming platform that handles the incoming data, streams gaze and worldview from multiple eye trackers, and streams the centralview from a network camera. Each data stream is sent to separate Kafka topics with multiple partitions. In Kafka, topics are analogous to sub-directories that store incoming streams to background Java programs called brokers, ensuring load balancing and scalability. Finally, separate consumer modules read these data streams for a) homography calculation, b) writing to local storage, or c) publishing the data to a dedicated Flask server for real-time monitoring.
@@ 3.1.3. Recording and streaming
3.2. Modules
@@ 3.2.1. GlassesRecord
The GlassesRecord module is engaged in the recording mode to send a recording start and stop signal to each pair of eye tracking glasses, controlled manually by an operator. Additionally, the module provides functionality to annotate recordings with custom hand-marked events. A critical challenge in multi-person eye tracking is the management and monitoring of many devices. An intuitive and reliable interface to this module is, therefore, a central requirement to efficiently trigger and monitor actions on multiple devices. We demonstrate an example implementation of this module in our Utility Test (see 4) using the open source Textual framework (https://textual.textualize.io/) to build a convenient Terminal User Interface (TUI). Further, the module ensures fault tolerance with the ability to trigger these actions selectively on a subset of devices, for example, the ability to restart recording on one pair of glasses if it fails, rather than having to restart on all devices.
Time-synchronisation. A critical aspect of working with multiple devices on a network is time synchronisation, since each device has a separate internal clock used to record timestamps on that device. Independent internal clocks must be synchronised to a master clock to ensure accurate timestamp comparisons. We synchronise all eye tracking devices, recording cameras, and internal servers to a central NTP server hosted locally on the network. These internal digital clocks, however, can also accumulate small drifts over time leading to errors in offset calculation. In the streaming mode of operation, this drift can be ignored for most applications to favour low-latency operation. In the recording mode of operation, the GlassesRecord module records periodic clock offset measurements–every ten seconds by default–between the NTP server clock and each eye tracking device clock. Offsets between the two clocks are measured by exchanging a brief sequence of UDP (User Datagram Protocol) packets between the two involved devices that yields a) an estimate of the round-trip-time (RTT) between the two devices, and b) an estimate of the clock offset (OFS) at the time of the measurement, with the round-trip-time factored out. The final offset estimate is the one offset value for which the associated RTT value was minimal. During offline analysis, the offset information for each device is aggregated over the entire span of the recording. An outlier-robust linear fit is calculated over the clock offset measurements using random sample consensus (RANSAC) and used to remap the timestamps for precise time synchronisation.

### v2

# SocialEyes: Scaling mobile eye-tracking to multi-person social settings

@@ Abstract.
Eye movements provide a window into human behaviour, attention, and interaction dynamics. Challenges in real-world, multi-person environments have, however, restrained eye-tracking research predominantly to single-person, in-lab settings. We developed a system to stream, record, and analyse synchronised data from multiple mobile eye-tracking devices during collective viewing experiences (e.g., concerts, films, lectures). We implemented lightweight operator interfaces for real-time-monitoring, remote-troubleshooting, and gaze-projection from individual egocentric perspectives to a common coordinate space for shared gaze analysis. We tested the system in a live concert and a film screening with 30 simultaneous viewers during each of two public events (N=60). We observe precise time-synchronisation between devices measured through recorded clock-offsets, and accurate gaze-projection in challenging dynamic scenes. Our novel analysis metrics and visualizations illustrate the potential of collective eye-tracking data for understanding collaborative behaviour and social interaction. This advancement promotes ecological validity in eye-tracking research and paves the way for innovative interactive tools.
multi-person, eye-tracking, social settings, naturalistic, visualisation, joint gaze, music, concert, film, homography
$c$s: Human-centered computing Usability testingccs: Human-centered computing Collaborative interactionccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing Visualization toolkitsccs: Human-centered computing Empirical studies in ubiquitous and mobile computingccs: Applied computing Performing artsccs: Software and its engineering Empirical software validation
@@ 1. Introduction
Eye contact and shared gaze are critical in social interactions and in guiding collective/joint attention. Eye gaze performs a unique dual function of perception and signalling (Gobel et al., 2015), making it an inherent part of human social behaviour. Tracking when, where, and how people shift their gaze in social interactions provides insights into the dynamics of individual as well as shared attention allocation (Valtakari et al., 2021; Clark, 1985), the influence of social context and co-presence (Foulsham et al., 2010), the role of non-verbal cues in communication (Senju and Csibra, 2008), and affective mechanisms such as the level of interest and engagement (Rahal and Fiedler, 2019). Moreover, due to its unobtrusive and direct nature, there has been a long-standing interest in interactive applications of eye-tracking (Jacob, 1990). These applications range from using eye movements as a proxy for traditional computer inputs (Ware and Mikaelian, 1986; Cui et al., 2023) to the development of innovative and creative tools (Creed et al., 2020; Hornof, 2014; Vamvakousis and Ramirez, 2016). Eye movements, therefore, serve as an exceptional tool for cognition research and human-computer interaction applications.
@@
Tracking eye movements has mostly been limited to single-person settings under controlled environments due to the immobility of traditional eye-tracking hardware and the complexities of unrestricted multi-person data collection and analysis. With the recent development of affordable, mobile eye-tracking technology, it is increasingly possible to conduct eye-tracking “in the wild” (see e.g. (Fasold et al., 2021; Kulke and Hinrichs, 2021; Reitstätter et al., 2020; Saxena et al., 2023; Vandemoortele et al., 2018; Fridman et al., 2018)). For single-person use cases, mobile eye-tracking glasses are most commonly applied to collect data in unrestricted settings but a number of challenges remain unsolved for their adoption in multi-person settings. For instance, these glasses record individual gaze from the egocentric perspective of the wearer; data from all differing perspectives of wearers need to be aligned when conducting multi-person studies. To effectively study collective gaze dynamics in naturalistic social settings, it is crucial to develop automated ways to synchronise and analyse the eye-tracking data of multiple individuals.
@@
In this paper, we look into the ubiquitous and cross-cultural social setting of collective viewing experiences where multiple people share a common gaze goal, e.g., sporting events (Kredel et al., 2017), surgery operations (Tolvanen et al., 2022), school classrooms (Jarodzka et al., 2021), etc. Eye-tracking in such contexts holds the potential of studying social cognition over extended timescales and developing novel interactive tools that promote group coordination and collaboration. For instance, while we have learned much about the cognitive neuroscience of music via laboratory studies with single participants and short sound clips presented in the lab, there is a push in the field for studying musical engagement in naturalistic social contexts (Tervaniemi, 2023), as the types of experiences people report as deeply meaningful often involve social settings with music that evolves over long periods of time.
A schematic of our approach to solve the problem of multi-person eye-tracking in such collective viewing experiences is detailed in Figure 1. We imagine a scenario where multiple people observe a shared scene from different viewing angles. The eye movements of each viewer are recorded independently from their egocentric perspective (worldview) using commercial-grade eye-tracking glasses. However, to study collective gaze patterns, the wearer-specific gaze data needs to be synchronised and mapped onto a shared coordinate space (semantic gaze mapping). We record a centralview of the shared scene from a camera placed behind all observers. We employ Network Time Protocol (NTP) to temporally synchronise data from all devices. The synchronised gaze data from all viewers is then mapped to the shared coordinate space i.e. the centralview by finding robust feature matches between worldview–centralview image pairs and performing a projective transformation. Following our approach we design a software framework–SocialEyes. We implement and test SocialEyes during two public events in which we recorded eye data from 30 participants simultaneously. We detail our choices of eye-tracking hardware, the underlying infrastructure, and the implementation specifics of the framework. In validating our approach, we provide visualisation and analysis metrics for multi-person eye-tracking data. Overall, we make the following contributions in this paper:
@@
    Develop and validate a scalable multi-person eye-tracking system, with flexible modes of operation. Our system provides users the means to monitor incoming data in real-time and to remotely control all associated mobile eye-tracking devices in parallel. We tested our system with 30 mobile eye-tracking devices in parallel, limited only by the number of devices we own. Our system should easily handle more devices when operating in recording mode. This tool opens unprecedented avenues for batch data collection and real-time interactivity in social settings.
@@
    Solve the problem of multi-perspective gaze data, in the context of shared social scenes. Using our proposed homography approach, it is possible to project the gaze of all eye-tracking glasses wearers from egocentric coordinates to shared central coordinates. This technique revolutionises possibilities for automating analyses of social gaze dynamics in real-world settings.
@@
    Introduce and implement standard and new visualisation and analysis techniques for multi-person eye-tracking data. We provide solutions for visualising both egocentric and homography-transformed gaze data and heatmaps, with methods to visualise all individually or in parallel. The transformed gaze heatmaps offer particular promise in analysing group attention and perception. Similarly, analysis metrics like convex hull area may offer novel insights into collective gaze behaviour.
Figure 1. LEFT: Schematic of a shared scene depicting multiple people wearing mobile eye-tracking glasses gazing at a stage. A central camera at the back of the audience records the shared scene (centralview). MIDDLE: Egocentric worldviews of each glasses wearer need to be projected onto the centralview. Point correspondences between the egocentric and centralviews are used to calculate homography matrices that relate the two views. RIGHT: Using the computed homography matrices to remap the gaze coordinates, all individuals’ gaze points can be projected onto the shared scene (centralview).
@@ 2. Related Work
Previous work on multi-person eye-tracking has relied on smaller groups (predominantly dyads) and elaborate experimental setups that hinder natural behaviour (Bise et al., 2024; Fasold et al., 2021; Macdonald and Tatler, 2018; Kera et al., 2016; Jermann et al., 2012; Kawase, 2009; Yamada et al., 2014). Most studies present static scenes on a screen as a proxy to predict real-world eye-tracking behaviour which was identified as an inefficient approach by Foulsham and Kingstone (2017). Some studies have applied head pose as a proxy for gaze in multi-person setups (Stiefelhagen, 2002; Beyan et al., 2018); however, existing research also shows that head orientation is not always correlated with gaze in social-group interactions (Vrzakova et al., 2016).
@@
In addition to the experimental setup, analysis of multi-person eye-tracking data remains a significant hurdle in conducting large scale studies. Previous studies have heavily relied on manual techniques like incorporating gaze reports in surveys (Pennill and Timmers, 2017), mapping gaze to visual stimuli using proprietary closed-source manufacturer software, or hand-coding regions of interest (Benjamins et al., 2018; Geeves et al., 2014). These techniques are incredibly time-consuming, with poor scalability to large datasets. Further, they are not entirely reproducible, and are prone to human errors. A common strategy applied to simplify the problem is to calculate coarse gaze measures, such as discrete binary classification of gaze on regions of interest (Müller et al., 2018; Fasold et al., 2021). While more scalable, such methods provide limited-resolution data, thereby reducing the cost-effectiveness of experiments and limiting the insights that can be gained.
@@
More scalable and automated attempts at unrestricted eye-tracking are actively being made under the umbrella field of ‘pervasive eye-tracking’ (Kasneci, 2017) and motivated by advances in portable head-mounted eye trackers that have recently become more affordable and accurate. The increased availability of open-sourced and automated analysis tools (Panetta et al., 2019; Li et al., 2006) has boosted the adoption and potential applications of these mobile eye trackers, which are increasingly being employed in unrestricted natural environments with static scenes, such as art galleries (Reitstätter et al., 2020). However, to date, a synchronised multi-person eye-tracking study in dynamic real world social settings has not been done and consequently the analysis methods to evaluate such large-scale data are not readily available. To solve these challenges we design and implement SocialEyes.
3. SocialEyes: Design Concept
@@
SocialEyes is a software framework that enables synchronous operation and analysis of multiple eye-tracking data streams in unrestricted, real-world settings. We followed the Research through Design approach (Stappers and Giaccardi, 2017) to design a software framework that is independent of specific hardware, network, and storage components, allowing it to adapt to various infrastructure constraints. In our software framework, presented in Figure 2, incoming data consists of three parallel streams: i) gaze coordinates from each eye-tracking device ii) worldview video recorded from a head-mounted camera on each device, and iii) a centralview video of the shared scene, captured and streamed by a stationary camera mounted to the centre of the scene. Incoming data streams are handled by separate modules that engage based on the mode of operation. Modules in our framework abstract information and allow isolated development and updates of separate components. Below, we describe the different modes of operation.
@@ 3.1.1. Recording
The recording mode is designed for offline operation of the framework where data are collected and stored for post-hoc analysis. Such operation is desirable when real-time processing of incoming data is not required, e.g., in research and user studies. Splitting the two stages of data collection and processing reduces network bandwidth and computation requirements, thereby allowing complex analysis and long-duration recordings with limited resources.
@@ 3.1.2. Streaming
The streaming mode is designed for real-time applications where data are collected and processed concurrently. Processing data in real-time is desirable for interactive applications such as gaze-contingent interfaces. This mode of operation demands sophisticated hardware and software optimisations to achieve good performance. Streaming of data in this mode is facilitated by the open-source event streaming platform–Apache Kafka–providing high-throughput and low-latency. Data streams are processed by Kafka topics, each comprising of multiple partitions that store data in background Java programs called brokers while ensuring load balancing and scalability. Finally, separate consumer modules read these data streams for a) calculating homography matrix, b) writing to local storage, and/or c) publishing the data to a dedicated Flask server for real-time monitoring.
@@ 3.1.3. Recording and Streaming
3.2. Technical Design
3.2.1. Time-synchronisation
. A critical challenge in managing multiple devices on a network is time synchronization, as each device relies on its internal clock to timestamp data samples. Independent internal clocks must be synchronised to a master clock to ensure accurate timestamp comparisons. All eye-tracking devices, recording cameras, and internal servers are, therefore, synchronised to a master NTP server accessible on the local network. This synchronisation ensures negligible clock offsets between devices, however, internal clocks on these devices can still accumulate variable drifts over long duration recordings. In the streaming mode of operation, this drift can be ignored for most applications to favour low-latency operation. However, for applications requiring fine-grained time alignment (typically in recording mode), we apply the clock-filter algorithm in NTP (Mills, 1994) to collect offset and drift at periodic intervals. Offsets between a reference device–the server running SocialEyes software– and each remote device (e.g., eye tracking glasses) are measured by exchanging a brief sequence of User Datagram Protocol (UDP) packets between the two involved devices. Each exchange returns a) an estimate of the round-trip-time (RTT) for the packet, and b) an estimate of the clock offset between the two devices at the time of the measurement, with the round-trip-time factored out. The final offset estimate is the one offset value for which the associated RTT value was minimal. Finally to align the timestamps from a device, its offset information is aggregated over the entire span of the recording (in recording mode) or a fixed size window (in streaming mode). An outlier-robust linear fit is calculated over the aggregated offset measurements using random sample consensus (Fischler and Bolles, 1981) to map the timestamps with the reference device. In our design, offset recording is performed as a background process by the GlassesRecord and GlassesStream modules for the recording and streaming mode respectively. The final synchronisation and linear mapping is performed as an optional process in each of the consumer modules i.e. Analysis, Visualisation, and Homography. We detail the functions of these modules below.
@@ 3.2.2. GlassesRecord
The GlassesRecord module is engaged in the recording mode to send recording start and stop signals to each pair of eye-tracking glasses. The module features an intuitive and reliable operator interface to efficiently trigger and monitor actions on multiple devices simultaneously (see example implementation in 4). Actions can be triggered selectively on a subset of devices to ensure fault tolerance, for example, the ability to restart recording on one pair of glasses if it fails, rather than having to restart on all devices. The module also provides functionality to annotate recordings with custom hand-marked events.